OpenStudy (anonymous):

Six builders worked 8 hours a day for 10 day to complete a job if 8 builders worked 6 hours a day for 3 days the remaining work has to be completed by 6 builders in 8 day. How many hours did each builder have to work each day to complete the remaining of the job? HELP ME PLEASE!?!

OpenStudy (anonymous):

let each worker do x amount of work in each day. so total work=8*6*10*x=480x the work done=6*8*3*x=144x remaining work=336x let t hours they have to do daily. so in 8 days 6 builder will do=8*6*t*x 48tx=336x t=336/48=7 hrs